# Balancing Diversity and Risk in LLM Sampling: How to Select Your Method and Parameter for Open-Ended Text Generation

### v1

@@ Abstract
Sampling-based decoding strategies have been widely adopted for Large Language Models (LLMs) in numerous applications, which target a balance between diversity and quality via temperature tuning and tail truncation (e.g., top-k and top-p sampling). Considering the high dynamic range of the candidate next-tokens given different prefixes, recent studies propose to adaptively truncate the tail of LLM’s predicted distribution. Although improved results haven been reported with these methods on open-ended text generation tasks, the results are highly dependent on the curated truncation parameters and exemplar text. In this paper, we propose a systematic way to estimate the intrinsic capacity of a truncation sampling method by considering the trade-off between diversity and risk at each decoding step, based on our collected prefix tree which preserves the context of a full sentence. Our work provides a comprehensive comparison between existing truncation sampling methods, as well as their recommended parameters as a guideline for users.
@@
Yuxuan Zhou1,3, Margret Keuper1,2, Mario Fritz3 1University of Mannheim, Germany 2Max Planck Institute for Informatics, Saarland Informatics Campus, Germany 3CISPA Helmholz Center for Information Security, Germany zhouyuxuanyx@gmail.com, keuper@mpi-inf.mpg.de, fritz@cispa.de
@@ 1 Introduction
Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023); Team et al. (2023) have demonstrated incredible performance in a variety of applications, and the reliability of decoding strategies has become a critical concern, especially where diverse and coherent samples are desired. Previous works have revealed that likelihood-maximization such as beam search Fan et al. (2018); Holtzman et al. (2020); Welleck et al. (2020); Meister et al. (2022) produces degenerated text which contains repetitive loops and incoherent context, particularly in the open-ended tasks. Therefore, sampling-based decoding strategies, e.g., Top-p Holtzman et al. (2020) and Top-k sampling Radford et al. (2018); Fan et al. (2018), have been widely adopted. The balance between diversity and quality of the generated text could be adjusted by tuning the temperature and truncation position to some extend, but requires non-trivial trial and error.
@@
We reveal that there exist two underlying issues in the current evaluation setup, which might hinder the assessment of a sampling method’s practical significance in real-world applications:
@@
The improvement of one method against another may be simply due to a better tuned parameter for the targeted exemplar text: the performance of sampling methods is sensitive to their parameters, and parameter sweep is often operated on a extremely sparse grid due to the high computation cost. This is especially problematic considering the non-linear dependency between performance and parameters.
@@
Users are agnostic to the optimal parameters in real-world applications: Practically speaking, users often pick parameters based on their own need for the compromise between diversity and quality, after few number of tryouts. There exits no universal optimal hyperparamters in different scenarios and users are agnostic to the optimal hyerparameters for their own tasks.
@@
The above issues exactly indicate the need for an evaluation that allows for estimating the theoretical capacity of a sampling method, independent of hyperparamter tuning. Moreover, the second issue additionally highlights the need for identifying the sweet spots of existing sampling methods, which could serve as a general guideline for parameter selection when applying a method.
@@
Based on the above analysis, we propose a systematic way to assess the inherent adaptability of a sampling method in different contexts. First, we rearrange Wikipedia-English 111https://dumps.wikimedia.org/ data in the form of a word-level prefix tree structure, or the so-called Trie Fredkin (1960); Ghasemi et al. (2019). As shown in fig. 3, all possible words that appear after a given prefix in the dataset are collected together as the child nodes, and their preceding word is regarded as the parent node. Starting from "Begin of Sequence" and collecting the child nodes recursively, we are able to transform the full dataset into a single prefix tree. It is noteworthy that a n-gram Trie Jurafsky (2000) tends to produce overestimated data support size given a prefix Bengio et al. (2000), due to the lost of contextual information outside the truncation window, as shown in fig. 1. In a similar spirit to Ding et al. (2024), we intentionally construct the prefix tree with only sentence-starting n-grams to preserve the context of a full sentence. thus we refer to the collected data as Context-Preserving Trie (CP-Trie).
Since the child nodes of the ancestral prefix define the data support, the recall of a sampling method w.r.t. data support is able to be computed. Though the training datasets of modern LLMs are significantly larger than Wikipedia dataset, such a data support serve as a reasonable lower-bound, especially when only sentence-level context are considered for evaluation, see section 4.1 for more details.
@@
Given the context-preserving prefix tree, we are able to evaluate the theoretical capacity of a sampling method, by examining the amount of tokens within and out of the data support with varying truncation parameter values. As can be seen in fig. 2, the optimal truncation position vary drastically given different prefixes and Top-k sampling could be regarded as a baseline method with zero adaptability. Therefore, an effective adaptive truncation mechanism is supposed to better follow such a variation, so that improved diversity can be achieved without harming the quality.
In summary, the contributions of our paper are as follows:
@@
We establish an intrinsic evaluation benchmark based on the Context-Preserving Prefix tree data, which allows for estimating the theoretical capacity of different sampling decoding methods via our proposed diversity and stability metrics.
@@
We conduct a comprehensive comparison of existing sampling approaches on our proposed benchmark, which serve as a guideline for choosing the truncation sampling methods and their corresponding parameter selection in real-world applications.
@@ 2 Related Work
In this section, we summarize the recently proposed sampling-based decoding strategies and the widely adopted benchmarks as well as metrics for open-ended text generation.
@@ 2.1 Sampling-based Decoding Methods
Vanilla sampling suffers from the risk of obtaining incoherent tokens, thus truncation of the tail distribution has been heavily discussed to alleviate such an issue, e.g., Top-k (Radford et al. (2018); Fan et al. (2018)) and Top-p sampling (Holtzman et al. (2020)). However, a fixed k or p is problematic when considering the high dynamic range of next reasonable tokens, as pointed out in more recent studies on adaptive sampling methods: Mirostat Basu et al. (2021) is proposed based on Zipf statistics and the assumption of a steady perplexity during generation. Hewitt et al. (2022) introduce -sampling which dismisses the tokens with low probabilities in the tail of the predicted distribution based on absolute and relative thresholds. Locally Typical Sampling Meister et al. (2023) assumes the generated text to retain a similar entropy rate to that of human-generated text. Adaptive Decoding Zhu et al. (2024) proposes to keep the entropy of the truncated distribution close to the original entropy. Although these approaches have been empirically evaluated on open-ended text generation tasks with curated truncation parameters and exemplar text, there still lacks a comprehensive comparison of their adaptability in more general cases.
@@ 2.2 Evaluation of Sampling-based Decoding
Benchmarks The commonly adopted benchmarks include story generation with WritingPrompts dataset Fan et al. (2018), document continuation with WikiText-103 dataset Merity et al. (2017) and abstractive summarization on the CNN/DAILYMAIL dataset Nallapati et al. (2016). These benchmarks suffer from the problem of limited exemplar text, which fails to capture the diverse nature of human language.
Statistical metrics are mostly based on n-gram statistics and focus on a single aspect, such as Repetition Welleck et al. (2020), Diversity Meister et al. (2023), Semantic coherence Gao et al. (2021), Zipf’s coefficient Holtzman et al. (2020) (simple Unigram rank-frequency statistics) and Self-BLEU Zhu et al. (2018).
Exemplar-based metrics dominate the evaluation of sampling-based decoding methods. As observed by Fan et al. (2018); Holtzman et al. (2020), lower perplexity of the generated text doesn’t necessarily indicate better quality. And Holtzman et al. (2020) suggested that the perplexity of the generated text is supposed to be close to that of the human text. MAUVE Pillutla et al. (2021) takes the trade-off between precision and recall into account, by comparing the learnt distribution from a text generation model to the distribution of human-written text using divergence frontiers. A recent study Shi et al. (2024a) provides a comprehensive evaluation on a large collection of tasks, which are mostly based on exemplar-based metrics.
@@ 3.1 Problem Formulation
where denotes the allowed set comprising candidate next-tokens for a given prefix, and is the renormalization term.
Given the Context-Preserving Trie of a reference dataset, we could compute the estimate of the optimal allowed set as follows :
@@
Let be the allowed set after truncation given the prefix . The optimal allowed set corresponds to the allowed set with the minimum size, while covering the full data support. It is the solution to the following objective function:
@@ 3.2 Remaining Issues
We reveal three major issues in the evaluation of truncation sampling. We first summarize the problem of directly using probability as quality metric, then show the choice of truncation parameter has a significant impact on the evaluation. Finally, we reveal that a minor difference of Recall and Risk values may result in significant changes in diversity and quality of the generated text.
@@
Word frequencies are average statistics across various topics, and assuming the optimal probabilities or the optimal ranking of each reasonable next token is ill-posed.
Impact of Parameter Selection We highlight the complexity and biases in parameter selection: Top-k and Top-p have constant upper bounds, i.e., the vocabulary size and 1, respectively. In contrast, the upper bounds of -sampling and adaptive sampling are dependent on LLM’s predicted distribution, because they truncate the tail distribution based on the likelihood of tokens and the slope of Min-Max scaled entropy, respectively. The importance of identifying the effective ranges of such parameters is also reflected in the authors’ choice of numeral digit for their proposed parameters. For example, is set to in Zhu et al. (2024) and is chosen from , and etc in Hewitt et al. (2022). In comparison, the adopted values for top-p sampling are merely two digits after zero, such as . Our analysis also shows the significance of identifying the sweet spots of different sampling methods.
The Butterfly Effect Although the top few samples possess the most probability mass, we reveal that a minor change in the size of the allowed set at each decoding step could lead to major differences in the quality of the generated text, due to LLMs’ auto-regressive nature. For example, the overall probability of at least one bad token to appear at the position in a sequence of length increases rapidly as increases (due to the product operator):
Similarly, if only probability mass is assigned to additional tokens at each step, for a sequence of length , there will be chances of obtaining extra diverse samples.
@@ 4 Method
In this section, we derive our method for evaluating different sampling-based decoding strategies. To circumvent the reliability issue of the probabilities we merely check whether the predicted next-token is in or out of the data support.
@@ 4.1 Probability-Independent Metrics
To quantify the diversity and quality of a sampling method based on CP-Trie, we define the Recall and Risk of a sampling method regarding a given prefix below:
@@
is dependent on the parameter selection for truncation, e.g., k value in top-k sampling. When the allowed set is smaller than the optimal one after truncation, Recall is smaller than one and Risk is regarded as zero. With further increased size of the allowed set, Recall reaches one but Risk emerges. Since the sizes of reasonable sets vary drastically for different prefixes, it is not possible to always retain the optimal allowed set with a predefined parameter. In this case, we reveal that the adaptability w.r.t. the varying size of data support of a sampling method indeed determines its effectiveness in real-world application.
Note that we ignore the risk of obtaining bad samples within the optimal allowed set, because such type of risk is unsolvable by truncation and is rather determined by the inherent capacity of the trained LLMs. However, such risk is less severe comparing to that introduced by inappropriate truncation, since LLMs exhibit a significant capability in predicting the next token Touvron et al. (2023); Achiam et al. (2023); Jiang et al. (2023); Team et al. (2023) and most bad samples reside in the tail distribution.
More importantly, our evaluation doesn’t rely on the empirical probability, which is biased and inaccurate due to limited dataset size or context window size. However, the tokens which appear in the dataset could be confidently regarded as reasonable, regardless of their actual probabilities. In addition, considering that temperature could change the flatness of distribution arbitrarily, we use ratio of token counts instead of probability mass to make the evaluation independent on temperature tuning and exemplar text. For a more detailed discussion with supporting examples, please refer to section A.2.
4.2 Parameter-Independent Evaluation
To eliminate the huge impact of parameter tuning on fair evaluation, we define the final diversity metric Recall at a given Risk level as follows:
@@ 4.2 Parameter-Independent Evaluation
Analogously, a family of critical values such as can be easily defined.
From the definition, it can be seen that such a metric is no longer dependent on the selection of parameter, thus it reflects the genuine capacity of a sampling method regardless of parameter tuning. This allows for a fair comparison between different sampling methods, especially considering their drastically different effective ranges, as mentioned in section 1 and section 3.2.
4.3 The Priority of Low Variance in Risk
Despite that more diverse text is desired for many generation tasks, the minimum requirement of the text to be coherent and reasonable is mostly prioritized in real-world applications. Besides the average recall at a given risk level, it is noteworthy that a stable adaptive truncation mechanism is also preferred , i.e., the variance of risks should be kept as low as possible at the given average risk level.
Conjecture 4.3.
At a given average risk level, the total amount of risk when generating a sequence of length is reduced with decreased variance of the risks at each decoding step.
The above conjecture can be mainly understood by the auto-regressive generation process of LLMs, as shown in eq. 3. The in-distribution probability dependent on the risk at each decoding step and is minimized when the product of the probabilities is maximized at a given average risk level. We could infer that the sum of the in-distribution probabilities is approximated determined at a given average risk level. For simplification, we assume their sum is unchanged. According to AM-GM inequality, the maximum of the product is achieved when each individual component is equal to each other. Roughly speaking, the less variance of the probability masses at each step, the larger their product and thus the smaller the total risk is.
@@ 5 Experiment
In this section, we conduct evaluation of existing sampling-based decoding approaches on our re-collected EnWiki CP-Trie dataset. We aim to estimate the inherent adaptability of exiting sampling-based methods and the results could be used as references for the application of LLMs in open-ended tasks.
@@ 5.1 Data Collection
As mentioned above, the construction of the CP-Trie is a re-collection of an existing dataset. We apply the described procedure to the English subset of Wikipedia dataset and name the resulting dataset EnWiki CP-Trie. Although the core idea is straight-forward to understand, we elaborate the main design choices in the following:
Basic Unit There are many possible units for splitting the datasets into individual fragments, such as article, paragraph, sentence and n-grams. Constructing a tree based on articles or paragraphs may require a larger amount of data than the training data of LLMs to guarantee an adequate number of branches (because LLMs lean to interpolate), whereas the construction based on n-grams suffers from poor contextual information and are heavily biased towards common tuplets of n tokens regardless of the context. Therefore, we adopt sentence as the basic unit for our dataset, which guarantees a coherent context at sentence-level and requires a much smaller amount of data than the training data.
Filtering To avoid invalid words or rare proper names which are unreasonable for the model to predict, we exclude the sentences containing such words by checking their presence in the WORD LIST dataset, which is available on the website 222word-list dataset homepage. It contains a total amount of 354986 words and explicitly excludes proper names and compound words. Section titles are also excluded, because they are often incomplete sentences with poor contextual information.
Statistics Wikipedia-English dataset has a total number of articles, which results in EnWiki CP-Trie with number of leaves after conversion, as shown in fig. 4.
Storage The extracted prefix tree is implemented as a nested dictionary and saved in a single JSON file. Since each lookup at any depth has constant complexity, the retrieval from our dataset is highly efficient. Moreover, the dictionary is easily extendable if extra data are needed for a more accurate estimation of the full data support.

### v2

@@ Abstract
Sampling-based decoding strategies have been widely adopted for Large Language Models (LLMs) in numerous applications, targeting a balance between diversity and quality via temperature tuning and tail truncation. Considering the strong dependency of the candidate next tokens on different prefixes, recent studies propose to adaptively truncate the tail of LLMs’ predicted distribution. Although improved results have been reported with these methods on open-ended text generation tasks, the results are highly dependent on the curated parameters and the limited exemplar text. In this paper, we propose a systematic way to estimate the capacity of a truncation sampling method by considering the trade-off between diversity and risk at each decoding step, based on our collected prefix tree which preserves the context of a full sentence. Our work offers a comprehensive comparison of existing truncation sampling methods and serves as a practical user guideline for their parameter selection. Our code is available at github repository.
@@
Yuxuan Zhou1,2, Margret Keuper1, Mario Fritz2 1University of Mannheim, Germany 2CISPA Helmholtz Center for Information Security, Germany yuxuan.zhou@cispa.de, keuper@uni-mannheim.de, fritz@cispa.de
@@ 1 Introduction
Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023); Team et al. (2023) have demonstrated exceptional performance across a variety of applications, and the reliability of decoding strategies has become a critical concern. Previous works have revealed that likelihood-maximization such as beam search Fan et al. (2018); Holtzman et al. (2020); Welleck et al. (2020); Meister et al. (2022) produces degenerate text which contains repetitive loops and incoherent context, particularly in open-ended tasks. Therefore, sampling-based decoding strategies, e.g., Top-p Holtzman et al. (2020) and Top-k sampling Radford et al. (2018); Fan et al. (2018), have been widely adopted. The balance between diversity and quality of the generated text could be adjusted by tuning the temperature and truncation position to some extend, but requires non-trivial trial and error.
@@
We reveal two underlying issues in the current evaluation, which hinder the assessment of a method’s significance in real-world applications:
@@
The improvement of one method over another might be simply due to a better tuned parameter for the targeted task: the performance of sampling methods is sensitive to their parameters, and parameter sweep is often operated on a extremely sparse grid due to the high computation cost. This is especially problematic considering the non-linear dependency between performance and parameters.
@@
Users are agnostic to the optimal parameters in real-world applications: Practically speaking, users often pick parameters based on their own need for the compromise between diversity and quality, after a few tryouts. There exists no universal optimal paramters in different scenarios and users are agnostic to the optimal parameters for their own tasks.
@@
The above issues exactly indicate the need for an evaluation that allows for estimating the theoretical capacity of a truncation sampling method (how well it adapts to the variation in data supports given different prefixes), independent of hyperparamter tuning. Moreover, the second issue additionally highlights the need to identify the sweet spots of existing sampling methods, which could serve as a user guideline for practitioners.
@@
In light of the above analysis, we propose a systematic way to assess the inherent adaptability of a sampling method. First, we rearrange Wikipedia-English 111https://dumps.wikimedia.org/ data into a word-level prefix tree, known as a Trie Fredkin (1960); Ghasemi et al. (2019). It is noteworthy that a n-gram Trie Jurafsky (2000) tends to overestimate the data support size given a prefix Bengio et al. (2000), as shown in Figure 1. In a similar spirit to Ding et al. (2024), we construct the prefix tree with only sentence-starting n-grams to preserve full sentence context, called Context-Preserving Trie (CP-Trie).
@@
Given the CP-Trie, we are able to estimate the theoretical capacity of a sampling method, by examining the amount of tokens within and out of the data support with varying truncation parameter values. As shown in Figure 2, the truncation positions, which exactly cover the full data supports, vary drastically given different prefixes and Top-k sampling could be regarded as a baseline method with zero adaptability. Therefore, an adaptive truncation method is supposed to better follow such a variation, so that improved diversity can be achieved without harming the quality.
In summary, our contributions are as follows:
@@
We establish an intrinsic evaluation benchmark based on the collected CP-Trie, which allows for estimating the theoretical capacity of different sampling methods via thoroughly designed diversity and stability metrics.
@@
We conduct a comprehensive comparison of existing sampling approaches, which serves as a guideline for choosing a method and its parameter in real-world applications.
We reveal that sampling-based decoding methods are underestimated in the existing study (Shi et al., 2024a) due to the difficulty in parameter selection, highlighting the merit of our evaluation protocol.
@@ 2 Related Work
In this section, we summarize recent sampling decoding strategies, along with common benchmarks and metrics for open-ended text generation.
@@ 2.1 Sampling-based Decoding Methods
Vanilla sampling suffers from the risk of obtaining incoherent tokens; thus, truncation of the tail distribution has been heavily discussed, e.g., Top-k (Radford et al., 2018; Fan et al., 2018) and Top-p sampling (Holtzman et al., 2020). However, a fixed k or p is problematic when considering the high dynamic range of next reasonable tokens, as pointed out in more recent studies on adaptive sampling methods: Mirostat Basu et al. (2021) is proposed based on Zipf statistics and the assumption of a steady perplexity during generation. Hewitt et al. (2022) introduce -sampling which dismisses the tokens with low probabilities in the tail of the predicted distribution based on absolute and relative thresholds. Locally Typical Sampling Meister et al. (2023) assumes that the generated text should retain a similar entropy rate to that of human-generated text. Adaptive Decoding Zhu et al. (2024) proposes to keep the entropy of the truncated distribution close to the original entropy. Although these approaches have been demonstrated to be effective, their performance is highly dependent on the curated truncation parameters and the limited exemplar text.
@@ 2.2 Evaluation of Sampling-based Decoding
Common benchmarks include story generation with WritingPrompts dataset Fan et al. (2018), document continuation with WikiText-103 dataset Merity et al. (2017) and abstractive summarization on the CNN/DAILYMAIL dataset Nallapati et al. (2016). These benchmarks suffer from the problem of limited exemplar text, which fails to capture the diverse nature of human language.
Statistical metrics are mostly based on n-gram statistics and focus on a single aspect, such as Repetition Welleck et al. (2020), Diversity Meister et al. (2023), Semantic coherence Gao et al. (2021), Zipf’s coefficient Holtzman et al. (2020) (Unigram rank-frequency) and Self-BLEU Zhu et al. (2018).
Exemplar-based metrics dominate the evaluation of sampling-based decoding methods. As observed by Fan et al. (2018); Holtzman et al. (2020), lower perplexity of the generated text does not necessarily indicate better quality. And Holtzman et al. (2020) suggested that the perplexity of the generated text should be close to that of the human text. MAUVE Pillutla et al. (2021) takes the trade-off between precision and recall into account, by comparing the learnt distribution from a text generation model to the distribution of human-written text using divergence frontiers. Shi et al. (2024a) provides a comprehensive evaluation on a large collection of tasks, mostly relying on exemplar-based metrics. However, we reveal that such evaluation is affected by the biases in the curated parameters and limited exemplar text, and our evaluation method is shown to alleviate such an issue.
@@ 3 Revisiting Truncation Sampling
We begin by revisiting the formulation of truncation sampling, followed by identifying the unresolved challenges in evaluating truncation sampling methods.
@@ 3.1 Problem Formulation
where denotes the allowed set of candidate next tokens at the position, given a sequence of tokens as prefix. is the renormalization term.
Given the Context-Preserving Trie of a reference dataset, we can compute the estimate of the optimal allowed set as follows :
@@
Let be the allowed set after truncation given the prefix . The approximated optimal allowed set corresponds to the allowed set with the minimum size, while covering the full data support for the token based on the Trie. It is the solution to the following objective function:
Note that the above definition is designed to exclude the risk of obtaining OOD tokens before the cutoff Finlayson et al. (2024), because such type of risk is unsolvable by truncation and is rather determined by the capacity of the trained LLMs. However, such risk is less severe compared to that introduced by inappropriate truncation, since LLMs exhibit a significant capability in predicting the next token Touvron et al. (2023); Achiam et al. (2023); Jiang et al. (2023); Team et al. (2023) and most OOD samples reside in the tail distribution.
@@ 3.2 Remaining Issues
We reveal three major issues in the evaluation of truncation sampling. We first summarize the problem of directly using probability as quality metric, then show that the choice of truncation parameter has a significant impact on the evaluation.
@@
Word frequencies are average statistics across various topics, and the optimal probabilities or ranking of each next token is ill-posed.
Parameter Sensitivity We highlight the complexity and biases in parameter selection: Top-k and Top-p have constant upper bounds, i.e., the vocabulary size and 1, respectively. In contrast, the upper bounds of -sampling and adaptive sampling are dependent on LLM’s predicted distribution, because they truncate the tail distribution based on the likelihood of tokens and the slope of Min-Max scaled entropy, respectively. The importance of identifying the effective ranges of such parameters is also reflected in the authors’ choice of numeral digit for their parameters. For example, is set to in Zhu et al. (2024) and is chosen from , and etc in Hewitt et al. (2022). In comparison, the adopted values for Top-p sampling are merely two digits after zero, such as . This shows the significance of identifying the sweet spots of different sampling methods.
@@ 4 Method
In this section, we derive our metrics for evaluating different sampling-based decoding strategies. The metrics are carefully designed to address the issues discussed in Section 3.2.
@@ 4.1 Probability-Independent Metrics
To circumvent the unreliable probability issue, we merely check whether the predicted next token is in or out of the data support. Specifically, we define Recall and Risk to quantify diversity and quality of a sampling method on a single node of CP-Trie:
@@
is dependent on the parameter selection for truncation, e.g., k value in Top-k sampling. When the allowed set is smaller than the approximated optimal allowed set after truncation, Recall is smaller than one and Risk is regarded as zero. With further increased size of the allowed set, Recall reaches one but Risk emerges. Since the sizes of reasonable sets vary drastically for different prefixes, it is not possible to always retain the approximated optimal allowed set with a predefined parameter. In this case, we reveal that the adaptability w.r.t. the varying size of data support of a sampling method indeed determines its effectiveness in real-world application.
More importantly, our evaluation does not rely on the empirical probability, which is biased and inaccurate due to limited dataset size or context window size. However, the tokens which appear in the dataset could be confidently regarded as reasonable, regardless of their actual probabilities. In addition, considering that temperature could change the flatness of distribution arbitrarily, we adopt ratio of token counts instead of probability mass to make the evaluation independent of temperature tuning and exemplar text. For a detailed discussion with supporting examples, please refer to Section A.2.
4.2 Tuning-Independent Evaluation
To eliminate the huge impact of Parameter Sensitivity issue on fair evaluation, we adopt Average Recall (AR) at an average Risk and Risk Standard Error (RSE) at an average Risk to quantify diversity and stability of a sampling method across nodes of CP-Trie, respectively:
@@ 4.2 Tuning-Independent Evaluation
where the superscript denotes the node in the evaluation set of nodes on the prefix tree. Analogously, a family of critical values such as can be easily defined.
Since is now determined by the given average Risk, the diversity metric reflects the genuine capacity of a sampling method regardless of parameter tuning. This allows for a fair comparison of different sampling methods, especially considering their drastically different effective ranges, as mentioned in Section 1 and Section 3.2.
@@ 5 Experiment
In this section, we conduct evaluation of existing sampling-based decoding approaches on our collected EnWiki CP-Trie dataset. We aim to estimate the inherent adaptability of sampling-based methods and the results could be used as references for the application of LLMs in open-ended tasks.
@@ 5.1 Data Collection
We construct our Trie data based on the English subset of Wikipedia dataset, named EnWiki CP-Trie. As shown in Figure 3, all possible words that appear after a given prefix in the dataset are treated as child nodes, with their preceding word regarded as the parent node. Starting from "Begin of Sequence" and collecting the child nodes recursively, we are able to transform the full dataset into a single prefix tree. We elaborate the main design choices in the following:
Basic Unit. It is possible to split the datasets into articles, paragraphs, sentences or n-grams. Constructing a tree based on articles or paragraphs may require more data than the training data of LLMs to guarantee an adequate number of branches (because LLMs lean to interpolate), whereas the construction based on n-grams suffers from poor contextual information and is heavily biased towards common tuplets of n tokens regardless of the context. Therefore, we adopt sentence as the basic unit, which guarantees a coherent context at sentence-level and requires much fewer data than training. It is noteworthy that a n-gram Trie Jurafsky (2000) tends to overestimate the data support size given a prefix Bengio et al. (2000), due to the loss of information outside the contextual window, as shown in Figure 1.
Filtering. To avoid invalid words or rare proper names which are unreasonable for the model to predict, we exclude the sentences containing such words by checking their presence in the WORD LIST dataset, which is available on the website 222word-list dataset homepage. It contains 354986 words in total and explicitly excludes proper names and compound words. Section titles are also excluded, which are often incomplete sentences with poor contextual information.
Statistics. Wikipedia-English dataset contains articles, which result in EnWiki CP-Trie with leaves, see Figure 4.
Storage. The prefix tree is implemented as a nested dictionary and saved in JSON format. Since each lookup at any depth has constant complexity, the retrieval is highly efficient. Moreover, the dictionary is easily extendable if extra data are needed for a more accurate estimation of the full data support.